You plan to invest in securities that pay 11.6%, compounded annually. If you invest $5,000 today, how many

    years will it take for your investment to grow to $9,140.20?

a.

6.10

b.

4.73

c.

6.43

d.

4.18

e.

5.50
Answer #1

We use the formula:
A=P(1+r/100)^n
where
A=future value
P=present value
r=rate of interest
n=time period.

9,140.20=5,000*(1.116)^n

(9,140.20/5,000)=(1.116)^n

Taking log on both sides;

log (9,140.20/5,000)=n*log (1.116)

n=log (9,140.20/5,000)/log (1.116)

=5.50 years(Approx).